Ages (years)

Frequency

Relative Frequency

15 - 19

20 - 24

25 - 29

30 - 34

35 - 39

40 - 44

45 - 49

Question 1 (1 point)

Please match the class with the correct frequency. You will need to determine the frequencies using the project data set and the variable "Age".

Question 1 options:

0

4

5

26

2

13

1.

class 15 - 19

2.

class 20 - 24

3.

class 25 - 29

4.

class 30 - 34

5.

class 35 - 39

6.

class 40 - 44
Answer #1

Solution

Ages (years Frequency Relative Frequency
15 - 19 0 0
20 - 24 4 4 / 50 =0.08
25 - 29 5 5 / 50 = 0.10
30 - 34 26 26 / 50 =0.52
35 - 39 2 2 / 50 =0.04
40 - 44 13 13 / 50 =0.26
45 - 49 0 0 / 50 =0
Total 50 1